# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

### v2

@@ 4.3 RNA-CM: Dynamic Function Modulation
(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNA (Roundtree et al., 2017). We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
(Processing): To build this task, we start from the entire dataset, partition it into connected components and apply our size filter. Then, we filter for systems that include modified residues, relying on PDB annotations. We apply our default redundancy removal and splitting strategies, which results in 185 data points.
Table 1: Description of the ML characteristics of each of our tasks. The number in parenthesis refers to the number of classes. We include published reference datasets available through rnaglib below their corresponding tasks.
Task name Level Type Dataset split sizes
1. RNA-GO RNA Multi-class (5) 349-75-75
2. RNA-IF Residue Multi-class (4) 1700-448-581
   gRNAde (Joshi et al., 2024) 11183-528-235
3. RNA-CM Residue Binary 138-29-30
4. RNA-Prot Residue Binary 881-189-189
5. RNA-Site Residue Binary 157-34-33
   RNASite (Su et al., 2021) 53-6-17
6. RNA-Ligand Pocket Multi-class (3) 203-43-44
7. RNA-VS (Carvajal-Patiño et al., 2025) Pocket Regression 304-34-65
@@ 4.4 RNA-Prot: Biological Complex Modeling
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a protein residue is closer than 8Å to any of its residues.
(Context): RNAs and proteins often bind to form a functional complex. Such complexes are involved in crucial cell processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008). RNA structure is often, but not always, heavily disrupted upon interaction with a protein.
(Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters. We remove systems originating from ribosomes that are fully encapsulated in a complex. Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, yielding 1251 data points.
@@
RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019). Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient. In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020). The following tasks aim to address questions in RNA small-molecule drug-design.
@@ 4.5 RNA-Site: Drug Target Detection
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a ligand is closer than 6Å to any of its residues.
(Context): The classical flow of a structure-based drug discovery pipeline starts with the identification of relevant binding sites, which are parts of the structure likely to interact with ligands, or of particular interest for a phenotypical effect. The structure of the binding site can then be used to condition the quest for small molecule binders, for instance using molecular docking (Ruiz-Carmona et al., 2014). The framing of this problem as a machine learning task for RNA was introduced in Rbind (Wang et al., 2018).
(Processing): We provide datasets and predefined splits from RNASite (Su et al., 2021), which are also utilized by other tools in the field such as RLBind (Wang et al., 2023). This dataset contains 76 systems obtained after applying stringent clustering on an older version of the PDB.
In addition to providing the existing dataset existing in the literature, we propose a larger, up-to-date dataset. We start by following similar steps as for RNA-Prot (without removal of ribosomal systems). We then include a drug-like filter on the small-molecules. Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only. We use the default redundancy removal and splitting, resulting in 458 systems.
@@ 4.6 RNA-Ligand: Pocket Categorization
(Definition): This is a binding site-level, multi-class classification task where the structure of an RNA binding site is classified according to the partner it accommodates.
(Context): Equipped with a binding site, one wants to use its structure to characterize its potential binders. On proteins, the Masif-Ligand tasks (Gainza et al., 2020) gathers all binding sites bound to the seven most frequent co-factors, and aims to classify them based on their partner. Inspired by this work we propose the RNA-Ligand task. To keep sufficiently many examples per task, we only retained the three most frequent classes. This task is less ambitious than training a molecular docking surrogate and can help understanding the potential modulators of a given binding site.
(Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules. We then proceed to extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder. This results in all binding pockets in our database and we group these pockets with a sequence clustering.
Then, our aim is to find the most frequent ligands binding in non-redundant pockets. To that end, we gather a dataset of binding site clusters that bind to one ligand only. We then compute the most frequent ligands among this dataset and retain only the top three. Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset and split based on structural similarity.
@@ 4.7 RNA-VS: Drug Screening
(Definition): This is a binding site-level, regression task. Given the structure of an RNA binding site and a small molecule represented as a molecular graph, the goal is to predict the affinity of their binding.
(Context): Beyond classification into fixed categories, virtual screening aims to score compounds based on their affinity to a binding pocket. This task is ubiquitous in drug design as it helps selecting the most promising compounds to be further assayed in the wet-lab. Our last task implements a virtual screening task introduced in (Carvajal-Patiño et al., 2025). Their model is trained to approximate normalized molecular docking scores. Trained models can then be used to rank compounds by their binding likelihood to target sites, hence achieving virtual screening.
(Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper. The authors curated a list of binding sites in a similar fashion to the one described in the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019). All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners. Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs.
@@ 5 Implementation
Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks. In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion. Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch (end-to-end reproducibility).
@@
Due to the rapid advances in the field, we can expect that additional interesting challenges will arise in the near future, complementary to the seven tasks introduced here. Thanks to the modularity of our tool, additional tasks on RNA can be easily integrated in our framework for future releases. This is illustrated in the documentation at rnaglib.org.
@@ 6 Experiments and Results
We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling. We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library. However, we provide a baseline performance for each of our proposed tasks as a reference for future works and conduct some investigations highlighting rnaglib’s use in understanding and building RNA structure-function models. All reported results are conducted over three seeds and code is available from github.com/BorgwardtLab/rnaglib-experiments.
@@ 6.1 Baseline Performance
The baseline performance for each of our proposed tasks is achieved by training Relational Graph Convolutional Networks (RGCN) on graph representations of the tasks’ underlying datasets. We choose to use 2.5D graphs, whose nodes represent residues and edges represent both canonical and non-canonical interactions. This approach was previously shown to outperform a secondary structure based alternative in a drug-discovery task (Oliver et al., 2020). We report our results in Table 2 and additional details on the experimental setup, including hyperparameters, and further metrics can be found in the supplementary Section A.2. We observe that on all tasks, our machine learning models are able to capture a meaningful signal that outperforms naive approaches, always predicting the majority class, validating the relevance of our tasks for benchmarking purposes.
@@
Table 2: Baseline performance reported using our recommended representative metric for each task. We compare to always predicting the majority class.
Task Metric Majority Score
RNA-GO Jaccard 0.2 0.396 $\pm$ 0.023
RNA-IF Sequence Recovery 0.28 0.372 $\pm$ 0.005
RNA-CM Balanced Accuracy 0.5 0.633 $\pm$ 0.007
RNA-Prot Balanced Accuracy 0.5 0.556 $\pm$ 0.004
RNA-Site Balanced Accuracy 0.5 0.655 $\pm$ 0.012
RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.124
RNA-VS AuROC 0.5 0.759 $\pm$ 0.005
Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al. (2024) for the RNA-VS task. Performance metrics are provided in Supplementary Tables 4 and 5. Despite using a significantly simpler model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
6.2 Redundancy and Splitting
To showcase rnaglib’s use in choosing suitable splitting approaches, we compare the three main splitting approaches it provides and present results in Figure 3(a). As expected, training a model on randomly split data leads to inflated results in some cases. However, the difference is not clear cut 3(a). Since our tasks’ datasets already pass a redundancy removal step, we postulate that the reduced redundancy mitigates the effect of leakage. Indeed, forgoing the redundancy removal for an example task (RNA-Site) markedly increases the performance difference between similarity-based and random splitting approaches.
rnaglib permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability. In Figure 3(b), we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used. Our results show that more stringent splitting leads to a decreased, more realistic performance. The trade-off between leakage and learning signal can thereby be tweaked by practitioners. We point out that structure and sequence thresholds are not directly comparable, since their calculation differs. Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting. This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures. Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, using both maximizes the reliability of the results.
(a) Figure 3a: Redundancy removal has a dampening effect on data leakage across splitting strategies.
(b) Figure 3b: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
6.3 Model Benchmarking
We provide two case studies on selected tasks elucidating how rnaglib can assist in choosing modeling modalities. We investigate the impact of increasing the depth of our model, as well as the effect of using different representations. We focus on RNA-Site, RNA-CM, and RNA-Prot.
First, we assess the impact of varying the number of message passing layers to understand the effect of varying structural context size on model performance. We train models of various depths (number of layers ranging from two to six) on each task of these three tasks, tuning only the learning rate which we adjust for optimal convergence, keeping the remaining hyperparameters as in Supplementary Table 3. Results are presented in Figure 4(a). While using four layers generally results in a good performance, the best value varies from one task to another. For example, predicting protein binding sites (RNA-Prot) might require less structural context than other tasks.
[…]
(a) Figure 4a: RGCN depth
(b) Figure 4b: Structural detail in input RNAs
Next, we assess various types of graph representations for RNA structures. We contrast 2.5D RNA graphs, whose edges include both canonical and non-canonical interactions, rnaglib’s default representation, with 2D RNA graphs (i.e. canonical interactions only). Within the latter, we consider a 2D setting, where all canonical interactions are encoded as one edge type, and a 2D+ setting, where canonical Watson-Crick base-pairing and backbone covalent interactions are distinguished (the backbone’s directionality results in a total of three edge types). We compare these three representations in Figure 4(b) and observe that including non canonical interactions results in a performance increase across all tasks. These results are in accordance with those presented in Oliver et al. (2020).

### v3

@@ 4.3 RNA-CM: Dynamic Function Modulation
(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides have been found in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNAs (Roundtree et al., 2017). We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
(Processing): We start with the entire dataset, partition it into connected components and apply our size filter. Then, we filter for systems that include modified residues, relying on PDB annotations. We apply our default redundancy removal and splitting strategies, which results in 197 data points.
@@ 4.4 RNA-Prot: Biological Complex Modeling
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether each residue is within 8Å of a protein residue.
(Context): RNAs and proteins often bind to form a functional complex. Such complexes are involved in crucial cellular processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008). RNA structure is often, but not always, heavily disrupted upon interaction with a protein.
(Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters. We remove systems originating from ribosomes that are fully encapsulated within a complex. Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, yielding 1251 data points.
@@
RNA is increasingly recognized as a promising class of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019). The ability to target RNA would drastically increase the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient. RNAs represent a therapeutic avenue in pathologies where protein targets are absent, e.g. triple-negative breast cancer (Xu et al., 2020). The following tasks address questions in RNA small-molecule drug-design.
@@ 4.5 RNA-Site: Drug Target Detection
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether each residue is closer than 6Å to a ligand.
(Context): The classic structure-based drug-discovery pipeline starts with the identification of relevant binding sites, which are parts of the structure likely to interact with ligands, or of particular interest for a phenotypic effect. The binding site structure can then be used to condition the search for small molecule binders, for instance using molecular docking (Ruiz-Carmona et al., 2014). The framing of this problem as a machine-learning task for RNA was introduced in Rbind (Wang et al., 2018).
(Processing): We provide datasets and splits from RNASite (Su et al., 2021), which are also utilized by other tools such as RLBind (Wang et al., 2023). This dataset contains 76 systems obtained after applying stringent clustering on an earlier version of the PDB. In addition to the published dataset, we propose a larger, up-to-date dataset. We start by following similar steps as for RNA-Prot (without removal of ribosomal systems), then apply a drug-like filter on the small-molecules. Finally, we remove systems with more than 10 protein atom neighbors, ensuring that RNA modulated binding. The default redundancy removal and splitting results in 224 systems.
@@ 4.6 RNA-Ligand: Pocket Categorization
(Definition): This is a binding site-level, multi-class classification task where the structure of an RNA binding site is classified according to the ligand it binds.
(Context): Equipped with a binding site, one wants to use its structure to characterize its potential binders. On proteins, the Masif-Ligand tasks (Gainza et al., 2020) gathers all binding sites bound to the seven most frequent cofactors, and aims to classify them based on their ligands. Inspired by this work, we propose the RNA-Ligand task. To retain sufficiently many examples per task, we only retained the three most frequent classes. This task is less ambitious than training a molecular docking surrogate and can help understanding the potential modulators of a given binding site.
(Processing): Starting with steps similar to RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules. We then extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder. This results in binding pockets, which we group by sequence clustering. To find the most frequent ligands binding in non-redundant pockets, we discard binding sites that bind to several ligands and retain only binding events to the three most frequent ligands in the remaining set. We split this final set based on structural similarity.
@@ 4.7 RNA-VS: Drug Screening
(Definition): This is a binding site-level regression task. Given the structure of an RNA binding site and a small molecule represented as a molecular graph, the goal is to predict their binding affinity.
(Context): Beyond fixed-category classification, virtual screening aims to score compounds based on their affinity to a binding pocket. This task is ubiquitous in drug design as it helps select promising compounds to be further assayed in the wet-lab. Here, we implement the virtual screening task introduced in (Carvajal-Patiño et al., 2025). Their model is trained to approximate normalized molecular docking scores, and can then rank compounds by their binding likelihood to target sites.
(Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper. The authors curated a list of binding sites analogously to the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019). All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners. Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs.
@@ 5 Implementation
Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks. In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choose a representation (in this example a Pytorch Geometric graph) and directly use the data in a simple and reproducible fashion. Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch, ensuring end-to-end reproducibility.
@@
Due to the rapid advances in the field, we can expect that additional interesting challenges will arise in the near future, complementary to the seven tasks introduced here. Thanks to the modularity of our tool, additional tasks on RNA can be easily integrated in our framework for future releases. This is illustrated in the documentation at rnaglib.org
@@ 6 Experiments and Results
We explicitly limit the scope of this work to the creation of a useful deep-learning library for RNA 3D structure based modeling. We refrain from suggesting novel model architectures and instead aim to illustrate the practical applicability of our library. Still, we provide a baseline performance for each of our proposed tasks as a reference for future works and demonstrate our library’s use in understanding and building RNA structure-function models.
All reported experiments were conducted over three seeds, to report error bars in our plots. Implementation details are provided in Section F. The code to reproduce our model training and our plots is available from github.com/BorgwardtLab/rnaglib-experiments.
6.1 Data Analysis and Timing
We characterize our tasks by analyzing the distribution of RNA families within them, primarily using Rfam (Ontiveros-Palacios et al., 2025), the RNA equivalent of Pfam (Mistry et al., 2021). However, due to a significant number of unannotated samples in Rfam, we also use the Nucleic Acid Knowledgebase (NAKB) (Lawson et al., 2023), the intended successor to the Nucleic Acid Database (NDB) (Coimbatore Narayanan et al., 2014), which provides broader coverage and additional nucleic acid–specific annotations. Across all tasks, we find a high number of tRNAs and rRNAs, which mirrors the distribution of RNA structures in the Protein Data Bank (PDB). A complete breakdown of these family and size distributions is available in Appendix C. Despite incomplete annotations, our datasets showcase the variety of functional roles played by RNAs, and thereby illustrate the suitability of our tasks for training models generalizable across families and molecular sizes.
Training a model on a task can be computationally demanding and thus a potential bottleneck. Given the modest amount of RNA structural data, we show that training on our tasks is feasible with limited compute, making the suite accessible to a wide variety of users. Concrete training times and memory requirements for each task are reported in Appendix D.
6.2 Redundancy and Splitting
Biological data points are inherently correlated due to evolutionary history, topological similarities, and minor experimental variations that introduce redundancy. Consequently, this data is not independent and identically distributed (i.i.d.). This non-i.i.d. nature means that commonly used random data splits can be insufficient (Volkov et al., 2022), which has motivated the development of more advanced splitting strategies (Bernett et al., 2024).
The choice of splitting strategy depends on the desired type of generalization, from simple interpolation to the more challenging extrapolation across dissimilar proteins or RNAs. The latter is driven by the ambition to use machine-learning to approximate the universal physical laws that are believed to govern structure-function relationships (Nori et al., 2025). To this end, similarity-aware evaluation allows researchers to precisely tune the desired degree of generalization, bridging the gap between random and strict similarity-based splits (Zhang et al., 2025).
To showcase our library’s utility in choosing suitable splitting approaches, we compare its three main splitting approaches in Figure 3. As expected, random splitting often leads to inflated performance metrics. However, this effect was less pronounced on our datasets, which we attribute to our initial redundancy removal step mitigating data leakage. Indeed, forgoing the redundancy removal for three example tasks markedly increases the performance gap between similarity-based and random splitting approaches.
Figure 3: Redundancy removal has a dampening effect on data leakage across splitting strategies.
For each splitting strategy, our library permits continuous tuning of the similarity thresholds, depending on use case and data availability. The trade-off between leakage and learning signal can thereby be tweaked by practitioners. In Supplementary Figure 7, we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used. Again, a more stringent splitting leads to a decreased performance.
6.3 Model Benchmarking
We provide two case studies on selected tasks elucidating how our library can assist in choosing modeling modalities. We investigate the impact of increasing the depth of our model, as well as the effect of using different RNA representations. We focus on RNA-Site, RNA-CM, and RNA-Prot.
We evaluated three classes of RNA representations reflecting its multi-level structure. 1D sequence models, such as LSTM or Transformers, use only the raw nucleotide sequence.
2D coarse-grained graphs represent nucleotides as nodes and encode structural information as edges. These range from simple base-pair and backbone connections considered as one (2D) or several edge types (2D+) to the 2.5D representation, which adds non-canonical interactions using Leontis-Westhof nomenclature (Leontis and Westhof, 2001). Those graphs are processed by a Relational Graph Convolutional Network (RGCN).
Finally, 3D geometric graphs explicitly model tertiary structure using residue coordinates, along with $k$-nearest neighbors edges (GVP) or with the 2.5D representation edges (GVP-2.5D). These graphs are processed by an equivariant graph neural network, the Geometric Vector Perceptron (GVP) Jing et al. (2020). Further details about the representations and models are provided in Appendix E.
[…]
(a) Figure 4a: Performance of different RNA representations: sequence (green), coarse-grained (blue) or full 3D (red).
(b) Figure 4b: Performance of the coarse grained approach for different RGCN depths.
Our results clearly indicate that coarse-grained models outperform 3D methods (Figure 4(a)). We hypothesize that atomic-level detail is unnecessary for modeling inherently unstable RNA structures, particularly with limited data. This is in line with existing findings on predicted structure Xu et al. (2024). However, our results also show that 3D methods outperform 1D methods, which contrast with the results on predicted structures, where noisy 3D information offered no advantage over sequence data. Our use of experimental structures demonstrates the definitive benefit of incorporating 3D structural information over 1D models.
Moreover, our results demonstrate that base-pairing information is a crucial prior for both 2D and 3D models, consistent with previous results obtained on a virtual screening task on graph representations Oliver et al. (2020). The 2.5D representation was the top performer, likely due to its combination of a powerful biological prior (non-canonical pairs) with an efficient, sparse graph structure.
Finally, we investigated the effect of structural context size by varying the depth of our 2.5D network from two to six layers (Figure 4(b)). While four layers consistently represented a strong baseline, the optimal depth is task-dependent. For instance, predicting protein binding sites (RNA-Prot) appears to require less structural context than other tasks.
@@ 6.4 Baseline Performance
Building on those results, we use our best-performing 2.5D approach for each of our proposed tasks to serve as a baseline. We report our results in Table 1 and additional details on the experimental setup, including hyperparameters, and further metrics are in the supplementary Section G. We also include the state-of-the-art performance for existing tasks. Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al. (2024) for the RNA-VS task. Performance metrics are in Supplementary Tables 8 and 9.
@@
Table 1: Baseline performance reported using our recommended representative metric for each task. We compare to always predicting the majority class. Scores with an asterisk refer to existing literature tasks and are reported directly from their respective publications.
Task Metric Majority Score Dataset split sizes
RNA-GO Jaccard 0.2 0.533 $\pm$ 0.013 349-75-75
RNA-IF Sequence Recovery 0.28 0.354 $\pm$ 0.005 1700-448-581
   gRNAde Sequence Recovery 0.28 0.568* 11183-528-235
RNA-CM Balanced Accuracy 0.5 0.661 $\pm$ 0.017 138-29-30
RNA-Prot Balanced Accuracy 0.5 0.553 $\pm$ 0.007 881-189-189
RNA-Site Balanced Accuracy 0.5 0.607 $\pm$ 0.005 157-34-33
   RNASite AuROC 0.5 0.703* 53-6-17
RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.006 203-43-44
RNA-VS AuROC 0.5 0.759 $\pm$ 0.005 304-34-65
Despite using a significantly simpler model, our approach achieves competitive results on those tasks, aligning with recent methods while falling short of state-of-the-art performance. We observe that on all tasks, our simple machine learning models are able to capture a meaningful signal that widely outperforms a naive approach of predicting the majority class. This confirms that our benchmark tasks are both learnable and challenging, providing a robust foundation for future model development.